OpenStudy (anonymous):

For the function y = f(x), what is the ordered pair for the point on the graph when x = 3n - 4? (3n - 4, f(n)) (3n - 4, f(3n - 4)) (x, f(3n)) (x, 3n - 4)

OpenStudy (anonymous):

I belive it is is the second one because it is given x = 3n-4 and y=f(x) =f(3n-4) So point P can be plotted by P(x,y) so i belive it is b i hope you understand better

OpenStudy (anonymous):

oh thanks for the medal haha

OpenStudy (anonymous):

no problem gorgeous :)